The wave function of the hydrogen atom in its ground state (1s) is = (
1
Expert's answer
2021-10-10T15:56:58-0400

The wave function of the hydrogen atom in its ground state (1s) is


"\psi_{100}=\frac1{\sqrt{\pi a_0^3}}e^{-\frac{r}{a_0}}."